2012-08-16find the total number of necleons present in 12g of12C atoms

Grade:12th Pass

1 Answers

Gaurav
9 years ago
12 g Carbon -12 = 1 mole
Atomic mass of c-12 = 24
Atomic number of c-12 = 12
Number of protons = 12
Number of neutrons = 24-12 = 12
Total number of nucleons in one C-12 atom = 12+12 = 24
Number of neucleons in 1 mole C-12 atoms = 6.023x 1023 x 24